# SORSA: Singular Values and Orthonormal Regularized Singular Vectors Adaptation of Large Language Models

### v1

@@ 5 Empirical Experiments
The results showed that on the Llama 2 7B model, SORSA achieved 10.36% accuracy on the MATH benchmark and 56.03% accuracy on the GSM-8K benchmark, both significantly outperforming PiSSA, LoRA, and Full FT. On the Mistral 7B v0.1 model, SORSA achieved 21.86% accuracy on the MATH benchmark and 78.03% accuracy on the GSM-8K benchmark, demonstrating slightly better performance on both tasks compared to the other three methods.
@@ 6 Conclusion
In this paper, we introduced SORSA, a novel parameter-efficient fine-tuning (PEFT) method designed to enhance the adaptation of large language models (LLMs) for downstream tasks. SORSA utilizes singular value decomposition (SVD) to split pre-trained weights into principle and residual components, only training the principle singular values and vectors while freezing the residuals. We implemented an orthonormal regularizer to maintain the orthonormality of singular vectors during training, ensuring efficient parameter updates and preserving the integrity of singular values.
@@
We adopt singular values and vectors analysis, comparing SORSA with partial FT and LoRA, and we find SORSA’s superiority in preserving the pre-trained weight’s singular values and vectors during training. This suggests an explanation for SORSA’s supreme performance demonstrated in the experiment. We also show the significance of the orthonormal regularizer through analysis. We also introduce the gradient analysis of SORSA in the paper, which demonstrates the process of scaling information to be transferred into $\Sigma_{p}$ during training, which makes the parameter updating more effective because it directly changes the global characteristics of the matrix.
@@ 7 Future Work
  • Quantized SORSA (QSORSA): Investigate the application of quantization techniques to SORSA introduced in QLoRA (Dettmers et al., 2024). This research aims to reduce the memory footprint and computational requirements of SORSA, making it feasible to deploy on edge devices or in environments with limited computational resources. By combining the benefits of quantization with the efficient fine-tuning capabilities of SORSA, QSORSA could enable highly efficient and powerful language models suitable for a broader range of practical applications.
@@ A.2 NLG Experiments
For our NLG tasks, we adapt Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023) models by LoRA (Hu et al., 2021), PiSSA (Meng et al., 2024) and SORSA, training with the first 100K data in MetaMathQA (Yu et al., 2024) dataset, and test by GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b) dataset, We use AdamW (Loshchilov & Hutter, 2018) optimizer and cosine annealing scheduler in training. SORSA layers are applied on all linear matrices in every layer. We only calculate the loss on the response part. The models are trained with BF16 mix precision. See Table 3 for hyperparameters.

### v2

@@ 5 Empirical Experiments
The results showed that on the Llama 2 7B model, SORSA achieved 10.36% accuracy on the MATH benchmark and 56.03% accuracy on the GSM-8K benchmark, both significantly outperforming PiSSA, LoRA, and Full FT. On the Mistral 7B v0.1 model, SORSA achieved 21.86% accuracy on the MATH benchmark and 78.03% accuracy on the GSM-8K benchmark, demonstrating slightly better performance on both evaluation tasks compared to the other three methods.
@@ 6 Conclusion
In this paper, we introduced SORSA, a novel parameter-efficient fine-tuning (PEFT) method designed to enhance the adaptation of large language models (LLMs) for downstream tasks. SORSA utilizes singular value decomposition (SVD) to split pre-trained weights into principal and residual components, only training the principal singular values and vectors while freezing the residuals. We implemented an orthonormal regularizer to maintain the orthonormality of singular vectors during training, ensuring efficient parameter updates and preserving the integrity of singular values.
@@
We adopt singular values and vectors analysis, comparing SORSA with partial FT and LoRA, and we find SORSA’s superiority in preserving the pre-trained weight’s singular values and vectors during training. This suggests an explanation for SORSA’s supreme performance demonstrated in the experiment. We also show the significance of the orthonormal regularizer through analysis. We also analyze the regularizer of SORSA in the paper, which demonstrates the process of scaling information to be transferred into $\Sigma_{p}$ during training, which makes the parameter updating more effective because it directly changes the global characteristics of the matrix.
@@ 7 Future Work
  • Quantization: Investigate the application of quantization techniques to SORSA similar to the approach of QLoRA (Dettmers et al., 2024). This research aims to reduce the memory footprint and computational requirements of SORSA, making it feasible to deploy on edge devices or in environments with limited computational resources. By combining the benefits of quantization with the efficient fine-tuning capabilities of SORSA, QSORSA could enable highly efficient and powerful language models suitable for a broader range of practical applications.
@@ A.2 NLG Experiments
For our NLG tasks, we adapt Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023) models by LoRA (Hu et al., 2021), PiSSA (Meng et al., 2024) and SORSA, training with the first 100K data in MetaMathQA (Yu et al., 2024) dataset, and test by GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b) dataset, We use AdamW (Loshchilov & Hutter, 2018) optimizer and cosine annealing scheduler in training. SORSA adapters are applied on all linear matrices in every layer. We only calculate the loss on the response part. The models are trained with BF16 mix precision. See Table 3 for hyperparameters.